OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)).

OpenStudy (anonymous):

-9x - 2 -9x + 5 -9x + 25 -9x + 27

OpenStudy (anonymous):

f(x) = 9x - 2 and g(x) = -x + 3 So f(g(x)) is simply plugging g(x) values in the f(x) sooo we write f(g(x)) as 9(-x + 3) - 2 -9x + 27 - 2 = -9x + 25